Vector = 6.0 m and points 30° south of east. Vector = 4.0 m and points 30° east of south. The resultant vector + is given by
1
Expert's answer
2020-10-25T18:29:28-0400

Solution

Magnitude of vector A= 6 m

Magnitude of vector B=4 m

Angle between A and B vectors

"\theta=30^°"

So resultant vector

"R=\sqrt{A^2+B^2+2AB\cos\theta}"

"R=\sqrt{6^2+4^2+2\times6\times4\cos 30^°}\\R=9.67m"